OpenStudy (anonymous):

Find the measure of the central angle that you would draw to represent 67% in a circle graph. Round your answer to the nearest degree.

OpenStudy (campbell_st):

this is a percentage of a quantity.... the quantity is 360 degrees so use \[\frac{percentage}{100} \times quantity\]

OpenStudy (anonymous):

so do i cross multiply or no?

OpenStudy (campbell_st):

nope no need to cross multiply just type it into a calculator as 67/100 x 360 =